OpenStudy (mariooandresss):

Solve 5 over x minus 5 equals the quantity of x over x minus 5, minus five fourths for x and determine if the solution is extraneous or not.x = -5, extraneous x = -5, non-extraneous x = 5, extraneous x = 5, non-extraneous

OpenStudy (mariooandresss):

\[\frac{ 5 }{ x-5 }=\frac{ x }{ x-5 }-\frac{ 5 }{ 4 }\]